### v1

# Investigating the Contextualised Word Embedding Dimensions Responsible for Contextual and Temporal Semantic Changes

Abstract
[…]
Investigating the Contextualised Word Embedding Dimensions Responsible for Contextual and Temporal Semantic Changes
Taichi Aida Tokyo Metropolitan University aida-taichi@ed.tmu.ac.jp Danushka Bollegala University of Liverpool danushka@liverpool.ac.uk
§ 1 Introduction
Meaning of a word is a dynamic phenomenon that is both contextual (i.e. depends on the context in which the word is used) Pilehvar and Camacho-Collados (2019) as well as temporal (i.e. the meaning of a word can change over time) Tahmasebi et al. (2021). A large body of methods have been proposed to represent the meaning of a word in a given context Devlin et al. (2019); Conneau et al. (2020); Zhou and Bollegala (2021), or within a given time period Hamilton et al. (2016); Rosenfeld and Erk (2018); Rosin and Radinsky (2022); Rosin et al. (2022); Aida and Bollegala (2023); Tang et al. (2023). In particular, sense-aware contextualised word embeddings such as XL-LEXEME Cassotti et al. (2023) obtained by fine-tuning MLMs such as XLM-RoBERTa Conneau et al. (2020) on WiC Pilehvar and Camacho-Collados (2019) have reported superior performance in SCD benchmarks, implying that semantic change related information can be accurately inferred from SCWEs.
Despite the above-reported empirical success, to the best of our knowledge, no prior work has investigated whether there are dedicated dimensions in the XL-LEXEME embedding space responsible for the semantic changes of the words it represents. In this paper, we study this problem from two complementary directions. First, in § 2, we investigate the embedding dimensions responsible for the contextual semantic changes of words using WiC as the evaluation task. Second, in § 3, we investigate the embedding dimensions responsible for the temporal semantic changes of words on SemEval-2020 Task 1 Schlechtweg et al. (2020) benchmark. In each setting, we compare pre-trained CWEs and the SCWEs obtained by fine-tuning on WiC using PCA and ICA, which have been used in prior work investigating dimensions in CWEs Yamagiwa et al. (2023). Our investigations reveal several interesting novel insights that will be useful when developing accurate and efficient low-dimensional SCD methods as follows.
PCA discovers contextual/temporal semantic change-aware axes within the top 10% of the transformed axes better than ICA.
In pre-trained embeddings, we identify a small number of axes responsible for contextual/temporal semantic changes, while such axes are uniformly distributed in the fine-tuned embeddings.
Using semantic change aware dimensions report comparable or superior performance over using all dimensions in SCD benchmarks.
§ 2 Contextual Semantic Changes
We first investigate the existence of axes responsible for contextual semantic changes. Recall that XL-LEXEME is fine-tuned from XLM-RoBERTa on WiC. Therefore, the emergence of any semantic change-aware axes due to fine-tuning can be investigated using contextual semantic change benchmarks. We use the test split of the English WiC dataset Pilehvar and Camacho-Collados (2019) for evaluations.
RQ: When do the contextual SCD-aware axes emerge?
To investigate whether contextual semantic change-aware axes were already present in the pre-trained CWEs, or do they emerge during the fine-tuning step, for each sentence-pair in WiC, we compute the difference between the two target word embeddings obtained from the pre-trained XLM-RoBERTa (CWEs) and the fine-tuned XL-LEXEME (SCWEs). If a particular axis is sensitive to contextual semantic changes, it will take similar values in the two target word embeddings, thus having a near-zero value in their subtraction. We then visualise the difference vectors for sentence pairs where the target word takes the same meaning in the two sentences (True) vs. different meanings (False).
As shown in Figure 1, we conduct this analysis for the non-transformed original axes (indicated as Raw here onwards), as well as for the PCA/ICA-transformed axes in order to investigate whether such transformations can discover the axes responsible for contextual semantic changes as proposed by Yamagiwa et al. (2023).222As in Yamagiwa et al. (2023), we used PCA and FastICA provided in scikit-learn https://scikit-learn.org/ . We see that the axes encoding contextual semantic changes are not obvious in the original CWEs after pre-training (Figure 1), but materialise during the fine-tuning process (Figure 1). Similar trends are observed with PCA-transformations (s 1 and 1), whereas ICA shows contrasting results (s 1 and 1). In contrast to prior recommendations for using ICA for analysing CWE spaces Yamagiwa et al. (2023), we find ICA to be less sensitive to contextual semantic changes of words.
RQ: Can top- PCA/ICA-transformed axes capture contextual semantic changes?
Yamagiwa et al. (2023) discovered that ICA-transformed axes represent specific concepts and their linear combinations could represent more complex concepts (e.g. cars italian ferrari). Based on this finding, we investigate whether a combination of top- axes can collectively represent contextual semantic changes of words. Specifically, we sort the axes in the descending order of their principal values and select the top- to represent a target word embedding. We then compute the Euclidean distance between CWEs of the target word in each sentence in each WiC test sentence-pair. We predict the target word to have the same meaning in the two sentences, if the Euclidean distance is below a threshold value. We vary this threshold and report Area Under the Curve (AUC) of Receiver Operating Characteristic (ROC) curves, where higher AUC values are desirable. In Figure 2, we show results for top of the PCA/ICA-transformed axes sorted by the explained variance ratio/skewness and compare against the baseline that uses all of the Raw dimensions.
For the pre-trained CWEs (Figure 2), we see that Raw reports slightly better AUC than PCA, but when fine-tuned (Figure 2) PCA matches Raw even by using less than 10% of the axes. On the other hand, ICA reports lower AUC values than both Raw and PCA in both models. These results indicate that PCA is better suited for discovering axes responsible for contextual semantic changes than ICA. We suspect that although ICA is able to retrieve concepts such as topics Yamagiwa et al. (2023), it is less fluent when discovering task-specific axes that require the consideration of different types of information. In conclusion, (1) contextual semantic change-aware axes emerge during fine-tuning, and (2) they are discovered by PCA even within 10% of the principal components.
§ 3 Temporal Semantic Changes
In contrast to contextual SCD, temporal SCD considers the problem of predicting whether a target word represents different meanings in two text corpora and , sampled at different points in time. For evaluations, we use the SemEval-2020 Task 1 English dataset Schlechtweg et al. (2020), which contains a manually rated set of target words for their temporal semantic changes. Further details on the tasks/datasets are given in Appendix A.
RQ: Can top- PCA/ICA-transformed axes capture temporal semantic changes?
Similar to Figure 2, we investigate whether PCA/ICA can discover axes responsible for temporal semantic changes by considering the top-% of axes for . We calculate the semantic change score of as the average pairwise Euclidean distance over the two sets of sentences containing the target word in and as conducted in previous work Kutuzov and Giulianelli (2020); Laicher et al. (2021); Cassotti et al. (2023). Finally, is predicted to have its meaning changed between and , if its semantic change score exceeds a pre-defined threshold. We vary this threshold and plot ROC in Figure 3.
In pre-trained CWEs, we can see that the use of the top 5% to 20% axes transformed by PCA is more effective in temporal semantic change detection than when all of the Raw dimensions are used (Figure 3). On the other hand, in fine-tuned SCWEs, Figure 3 indicates that PCA-transformed axes achieve the same AUC scores as Raw, similar to the contextual semantic change (Figure 2). Similar to the observation in contextual semantic change, ICA returns lower performance than both Raw and PCA.
To further investigate whether the top PCA/ICA axes can explain the degree of temporal semantic change, we measure the Spearman correlation between the semantic change scores and human ratings available in the SemEval-2020 Task 1 following the standard evaluation protocol for this task Rosin et al. (2022); Rosin and Radinsky (2022); Aida and Bollegala (2023); Cassotti et al. (2023); Periti et al. (2024). As shown in Figure 4 for the pre-trained CWEs (Figure 4), using only 10% of the axes, PCA outperforms Raw that uses all axes. Moreover, for the fine-tuned SCWEs (Figure 4), using only 10% of the axes PCA achieves the same performance as Raw. However, ICA consistently underperforms in both pre-trained and fine-tuned settings.
§ 4 Conclusion
We found that there exists a smaller number of axes that encode contextual and temporal semantic changes of words in MLMs, which are accurately discovered by PCA. These findings have several important practical implications. First, it shows that MLMs can be compressed into efficient and accurate lower-dimensional embeddings when used for SCD tasks. Second, it suggests the possibility to efficiently update a pre-trained MLM to capture novel semantic associations of words since the MLM was first trained, by updating only a smaller number of dimensions. Exploring these opportunities is beyond the scope of this short paper and is deferred to future work.
Limitations
To simplify our analysis, we limited our experiments reported in the paper to English, which is a morphologically limited language. Although our method of analysis is language agnostic, it is essential to repeat these experiments on datasets other than English to confirm the generalisability of the conclusions. For instance, the SemEval-2020 Task 1 dataset includes Latin, German, and Swedish datasets, in addition to English, and there are also WiC datasets available in multiple languages. Notably, our analysis only requires pre-trained or publicly available fine-tuned MLMs, which are available for all those languages, which makes such a multilingual analysis possible.
Ethical Considerations
In this paper, we focus on investigating the existence of dedicated dimensions about contextual/temporal semantic changes of words. For the best of our knowledge, no ethical issues have been reported for the WiC and SCD datasets we used in our experiments. On the other hand, we also used publicly available pre-trained/fine-tuned MLMs, some of which are known to encode and potentially amplify unfair social biases (Basta et al., 2019). Whether such social biases are influenced by the dimension selection methods we consider in the paper must be carefully evaluated before deploying any MLMs in downstream applications.
Acknowledgements
This work was supported by JST, the establishment of university fellowships towards the creation of science technology innovation, Grant Number JPMJFS2139.
Appendix A Examples of Contextual/Temporal Semantic Change Task
In this section, we further explain the two types of semantic changes of words considered in the paper: (a) contextual semantic changes and (b) temporal semantic change.
Contextual Semantic Change Detection Task involves predicting whether the meaning of a word in a given pair of sentences are the same. For example, an ambiguous word can express different meanings in different contexts, which is considered under contextual semantic changes. Models are required to make a prediction for each pair of sentences.
Temporal Semantic Change Detection Task involves predicting the meanings of a word in given sets of sentences across different time periods. A word that was used in a different meaning in the past can be associated with novel meanings later on, which is considered as a temporal semantic change of that word. Models predict whether the meaning of the word has changed over time by comparing the given sets of sentences.
Both types of semantic changes are common and even the same word can undergo both types of semantic changes as shown in Table 1 for the word plane. The contextual semantic change task requires models to be sensitive to the context within just two given sentences, whereas the temporal semantic change task requires models to account for the semantic changes of words across two different time periods.
| Type of Semantic Change | Instances | Label | |
| Contextual | …two points on a plane lies … | …the plane graph as the X-Y … | True (Same meanings) |
| He lived on a worldly plane. | …the plane graph as the X-Y … | False (Different meanings) | |
| Temporal | • …this is a horizontal plane, and … • …because it is parallel with the ground plane … • …this is a horizontal plane, … | • …as the plane settled down at … • …558 combat planes and 4,000 tanks. • The President’s plane landed at … | True (Semantically Changed) |