### v2

# Investigating the Contextualised Word Embedding Dimensions Specified for Contextual and Temporal Semantic Changes

Abstract
[…]
Investigating the Contextualised Word Embedding Dimensions Specified for Contextual and Temporal Semantic Changes
Taichi Aida Tokyo Metropolitan University aida-taichi@ed.tmu.ac.jp Danushka Bollegala University of Liverpool danushka@liverpool.ac.uk
§ 1 Introduction
Meaning of a word is a dynamic phenomenon that is both contextual (i.e. depends on the context in which the word is used) Pilehvar and Camacho-Collados (2019) as well as temporal (i.e. the meaning of a word can change over time) Tahmasebi et al. (2021). A large body of methods have been proposed to represent the meaning of a word in a given context Devlin et al. (2019); Conneau et al. (2020); Zhou and Bollegala (2021); Rachinskiy and Arefyev (2021); Periti et al. (2024), or within a given time period Hamilton et al. (2016); Rosenfeld and Erk (2018); Aida et al. (2021); Rosin et al. (2022); Aida and Bollegala (2023b); Tang et al. (2023); Fedorova et al. (2024). In particular, sense-aware contextualised word embeddings such as XL-LEXEME Cassotti et al. (2023) obtained by fine-tuning masked language models such as XLM-RoBERTa Conneau et al. (2020) on Word-in-Context (WiC) Pilehvar and Camacho-Collados (2019) have reported superior performance in SCD benchmarks Cassotti et al. (2023); Aida and Bollegala (2023a); Periti and Tahmasebi (2024); Aida and Bollegala (2024), implying that semantic changes can be accurately inferred from SCWEs.
Type of Semantic Change | Instances | Label | |
Contextual | …two points on a plane lies … | …the plane graph as the X-Y … | True (Same meanings) |
He lived on a worldly plane. | …the plane graph as the X-Y … | False (Different meanings) | |
Temporal | • …this is a horizontal plane, and … • …because it is parallel with the ground plane … • …this is a horizontal plane, … | • …as the plane settled down at … • …558 combat planes and 4,000 tanks. • The President’s plane landed at … | True (Semantically Changed) |
Despite the empirical success, to the best of our knowledge, no prior work has investigated whether there are dedicated dimensions in the XL-LEXEME embedding space specified for the semantic changes of the words it represents. In this paper, we study this problem from two complementary directions. First, in § 3, we investigate the embedding dimensions specific to the contextual semantic changes of words using WiC benchmarks Pilehvar and Camacho-Collados (2019); Raganato et al. (2020); Martelli et al. (2021); Liu et al. (2021) as the evaluation task. Second, in § 4, we investigate the embedding dimensions specific to the temporal semantic changes of words on SemEval-2020 Task 1 Schlechtweg et al. (2020) benchmark. In each setting, we compare pre-trained CWEs and the SCWEs obtained by fine-tuning on WiC using PCA and ICA, which have been used in prior work investigating dimensions in CWEs Yamagiwa et al. (2023). Our investigations reveal several interesting novel insights that will be useful when developing accurate and efficient low-dimensional SCD methods as follows.
PCA discovers contextual/temporal semantic change-aware axes within the top 10% of the transformed axes better than ICA.
In pre-trained embeddings, we identify a small number of axes that are specified for contextual/temporal semantic changes, while such axes are uniformly distributed in the fine-tuned embeddings.
Semantic change aware dimensions report comparable or superior performance over using all dimensions in SCD benchmarks.
§ 2 Task Description
In this section, we explain the two types of semantic changes of words considered in the paper: (a) contextual semantic changes and (b) temporal semantic changes.
Contextual Semantic Change Detection Task
involves predicting whether the meaning of a word in a given pair of sentences are the same Pilehvar and Camacho-Collados (2019). For example, an ambiguous word can express different meanings in different contexts, which is considered under contextual semantic changes. Models are required to make a prediction for each pair of sentences.
Temporal Semantic Change Detection Task
involves predicting the meanings of a word in given sets of sentences across different time periods Schlechtweg et al. (2020). A word that was used in a different meaning in the past can be associated with novel meanings later on, which is considered as a temporal semantic change of that word. Models predict whether the meaning of the word has changed over time by comparing the given sets of sentences.
Models
For the Contextual Semantic Change Detection Task, contextual word embeddings Devlin et al. (2019); Conneau et al. (2020) are the primary choice, as they effectively capture word meanings based on sentence context. For the Temporal Semantic Change Detection Task, both static Kim et al. (2014); Kulkarni et al. (2015); Hamilton et al. (2016); Yao et al. (2018); Aida et al. (2021) and contextual Rosenfeld and Erk (2018); Kutuzov and Giulianelli (2020); Laicher et al. (2021); Aida and Bollegala (2023b) embeddings can be applied. Notably, sense-aware contextual embeddings trained specifically for contextual semantic change tasks have achieved superior performance, demonstrating their broader applicability Cassotti et al. (2023); Aida and Bollegala (2024).
Both types of semantic changes are common and even the same word can undergo both types of semantic changes as shown in Table 1 for the word plane. The contextual semantic change task requires models to be sensitive to the context within just two given sentences, whereas the temporal semantic change task requires models to account for the semantic changes of words across two different time periods.
§ 3 Contextual Semantic Changes
We first investigate the existence of axes specific to contextual semantic changes. Recall that XL-LEXEME is fine-tuned from XLM-RoBERTa on WiC datasets. Therefore, the emergence of any semantic change-aware axes due to fine-tuning can be investigated using contextual semantic change benchmarks. We use the test split of the English WiC Pilehvar and Camacho-Collados (2019), XL-WiC (Raganato et al., 2020), MCL-WiC (Martelli et al., 2021), and AM2iCo (Liu et al., 2021) datasets for evaluations.222Due to the page limitations, results for other datasets than the English WiC are shown in Appendix B. Data statistics are in Appendix A.
RQ1: When do the contextual SCD-aware axes emerge?
To investigate whether contextual semantic change-aware axes were already present in the pre-trained CWEs, or do they emerge during the fine-tuning step, for each sentence-pair in WiC datasets, we compute the difference between the two target word embeddings obtained from the pre-trained XLM-RoBERTa (CWEs) and the fine-tuned XL-LEXEME (SCWEs). To obtain the sets of target word embeddings, we follow Cassotti et al. (2023) by using a Sentence-BERT Reimers and Gurevych (2019) architecture. We conduct this analysis for the non-transformed original axes (indicated as Raw here onwards), as well as for the PCA/ICA-transformed axes in order to investigate whether such transformations can discover the axes specified for contextual semantic changes as proposed by Yamagiwa et al. (2023).333As in Yamagiwa et al. (2023), we used PCA and FastICA provided in scikit-learn https://scikit-learn.org/ . In this paper, PCA/ICA-transformed axes are sorted by the experimental variance ratio/skewness, and this process is consistently applied where PCA or ICA is used. If a particular axis is sensitive to contextual semantic changes, it will take similar values in the two target word embeddings, thus having a near-zero value in their subtraction.
To address RQ1, we visualised the difference vectors for sentence pairs where the target word takes the same meaning in the two sentences (True) vs. different meanings (False). This visualisation was performed by following steps: (a) we prepared Raw or PCA/ICA-transformed axes; (b) for each WiC instance, which contains two sentences and a label, we calculated the difference between pair of sentences; (c) we normalised each axis (min=0 and max=1) for visualisation purposes.
As shown in Figure 1, we see that the axes encoding contextual semantic changes are not obvious in the original CWEs after pre-training (Figure 1), but materialise during the fine-tuning process (Figure 1). Similar trends are observed with PCA-transformations (s 1 and 1), whereas ICA shows contrasting results (s 1 and 1). In contrast to prior recommendations for using ICA for analysing CWE spaces Yamagiwa et al. (2023), we find ICA to be less sensitive to contextual semantic changes of words. Interestingly, similar results have been shown in other languages/datasets(Appendix B). 444Our findings do not aim to claim the superiority of PCA over ICA but to explore the existence of task-specific axes. Experimental results show that for semantic change tasks, PCA provides more task-related axes because (a) PCA orders axes by importance (eigenvalue), making task-related axes more accessible, and (b) ICA-transformed axes require external sorting method based on skewness rather than importance. Prior research indicates that ICA can capture topic-related axes Yamagiwa et al. (2023), suggesting that ICA may still hold potential for obtaining task-related axes. Further refinement of the approach remains as future research.
RQ2: Can top- PCA/ICA-transformed axes capture contextual semantic changes?
Yamagiwa et al. (2023) discovered that ICA-transformed axes represent specific concepts and their linear combinations could represent more complex concepts (e.g. cars italian ferrari). Based on this finding, we investigate whether a combination of top- axes can collectively represent contextual semantic changes of words. Specifically, we select the top- of the axes to represent a target word embedding. We then compute the Euclidean distance between CWEs of the target word in each sentence for every test sentence-pair in the WiC datasets. We predict the target word to have the same meaning in the two sentences, if the Euclidean distance is below a threshold value. We vary this threshold and report Area Under the Curve (AUC) of Receiver Operating Characteristic (ROC) curves, where higher AUC values are desirable. In Figure 2, we show results for top of the PCA/ICA-transformed axes and compare against the baseline that uses all of the Raw dimensions.
For the pre-trained CWEs (Figure 2), we see that Raw reports slightly better AUC than PCA, but when fine-tuned (Figure 2) PCA matches Raw even by using less than 10% of the axes. On the other hand, ICA reports lower AUC values than both Raw and PCA in both models. These results indicate that PCA is better suited for discovering axes specified for contextual semantic changes than ICA. We suspect that although ICA is able to retrieve concepts such as topics Yamagiwa et al. (2023), it is less fluent when discovering task-specific axes that require the consideration of different types of information. In conclusion, (1) contextual semantic change-aware axes emerge during fine-tuning, and (2) they are discovered by PCA even within 10% of the principal components. Notably, in other languages/datasets, similar trends have been observed (Appendix B). These results suggest that contextual semantic change-aware dimensions can be observed within 10% of the PCA-transformed axes across different languages.
§ 4 Temporal Semantic Changes
In contrast to contextual SCD, temporal SCD considers the problem of predicting whether a target word represents different meanings in two text corpora and , sampled at different points in time. For evaluations, we use the SemEval-2020 Task 1 dataset555Data statistics are in Appendix A. Schlechtweg et al. (2020), which contains a manually rated set of target words for their temporal semantic changes in English, German, Swedish, and Latin.666Due space limitations, results for languages other than English are shown in Appendix B.
RQ3: Can top- PCA/ICA-transformed axes capture temporal semantic changes?
Similar to Figure 2, we investigate whether PCA/ICA can discover axes specified for temporal semantic changes by considering the top-% of axes for . We calculate the semantic change score of as the average pairwise Euclidean distance over the two sets of sentences containing the target word in and as conducted in previous work Kutuzov and Giulianelli (2020); Laicher et al. (2021); Cassotti et al. (2023). Finally, is predicted to have its meaning changed between and , if its semantic change score exceeds a pre-defined threshold. We vary this threshold and plot ROC in Figure 3.
In pre-trained CWEs, we can see that the use of the top 5% to 20% axes transformed by PCA is more effective in temporal semantic change detection than when all of the Raw dimensions are used (Figure 3). On the other hand, in fine-tuned SCWEs, Figure 3 indicates that PCA-transformed axes achieve the same AUC scores as Raw, similar to the contextual semantic change (Figure 2). Similar to the observation in contextual semantic change, ICA returns the lowest performance.
To further investigate whether the top PCA/ICA axes can explain the degree of temporal semantic change, we measure the Spearman correlation between the semantic change scores and human ratings available in the SemEval-2020 Task 1 following the standard evaluation protocol for this task Rosin et al. (2022); Rosin and Radinsky (2022); Aida and Bollegala (2023b); Cassotti et al. (2023); Periti and Tahmasebi (2024); Aida and Bollegala (2024). As shown in Figure 4 for the pre-trained CWEs (Figure 4), using only 10% of the axes, PCA outperforms Raw that uses all axes. Moreover, for the fine-tuned SCWEs (Figure 4), using only 10% of the axes PCA achieves the same performance as Raw. However, ICA consistently underperforms in both pre-trained and fine-tuned settings. Importantly, we see similar trends in other languages (Appendix B). These results suggest that temporal semantic change-aware dimensions can also be observed within 10% of PCA-transformed axes across different languages.
§ 5 Conclusion
We found that there exists a smaller number of axes that encode contextual and temporal semantic changes of words in MLMs, which are accurately discovered by PCA. These findings have several important practical implications. First, it shows that MLMs can be compressed into efficient and accurate lower-dimensional embeddings when used for SCD tasks. Second, it suggests the possibility of efficiently updating a pre-trained MLM to capture novel semantic associations of words since the MLM was first trained, by updating only a smaller number of dimensions.
Limitations
In this paper, we limited experiments to XLM-RoBERTa based MLM models. These models are all fine-tuned on WiC datasets and have reported state-of-the-art (SoTA) performance in SCD benchmarks. We consider it would be important to further validate the findings reported in this paper using other embedding models and across multiple downstream applications.
Ethical Considerations
In this paper, we focus on investigating the existence of dedicated dimensions capturing contextual/temporal semantic changes of words. For the best of our knowledge, no ethical issues have been reported for the WiC and SCD datasets we used in our experiments. On the other hand, we also used publicly available pre-trained/fine-tuned MLMs, some of which are known to encode and potentially amplify unfair social biases (Basta et al., 2019). Whether such social biases are influenced by the dimension selection methods we consider in the paper must be carefully evaluated before deploying any MLMs in downstream applications.
Acknowledgements
Taichi Aida would like to acknowledge the support by JST, the establishment of university fellowships towards the creation of science technology innovation, Grant Number JPMJFS2139.
Appendix A Data Statistics
Full statistics of contextual and temporal SCD benchmarks are shown in Table 2 and Table 3.777WiC, XL-WiC, and MCL-WiC are licensed under the Creative Commons Attribution-NonCommercial 4.0 License, while AM2iCo and SemEval-2020 Task 1 are licensed under the Creative Commons Attribution 4.0 International License.
Dataset | Language | #Train | #Dev | #Test |
Monolingual | ||||
WiC | English | 5.4k | 6.4k | 1.4k |
XL-WiC | German | 48k | 8.9k | 1.1k |
French | 39k | 8.6k | 22k | |
Italian | 1.1k | 0.2k | 0.6k | |
MCL-WiC | Arabic | - | 0.5k | 0.5k |
English | 4.0k | 0.5k | 0.5k | |
French | - | 0.5k | 0.5k | |
Russian | - | 0.5k | 0.5k | |
Chinese | - | 0.5k | 0.5k | |
Cross-lingual | ||||
AM2iCo | German | 50k | 0.5k | 1.0k |
Russian | 28k | 0.5k | 1.0k | |
Japanese | 16k | 0.5k | 1.0k | |
Chinese | 13k | 0.5k | 1.0k | |
Arabic | 9.6k | 0.5k | 1.0k | |
Korean | 7.0k | 0.5k | 1.0k | |
Finnish | 6.3k | 0.5k | 1.0k | |
Turkish | 3.9k | 0.5k | 1.0k | |
Indonesian | 1.6k | 0.5k | 1.0k | |
Basque | 1.0k | 0.5k | 1.0k |
Language | Time Period | #Targets | #Tokens |
English | 1810–1860 | 37 | 6.5M |
1960–2010 | 6.7M | ||
German | 1800–1899 | 48 | 70.2M |
1946–1990 | 72.3M | ||
Swedish | 1790–1830 | 31 | 71.0M |
1895–1903 | 110.0M | ||
Latin | B.C. 200–0 | 40 | 1.7M |
0–2000 | 9.4M |
Appendix B Full Results
In this section, we present the full results of contextual and temporal SCD tasks. For the contextual SCD, visualisations of instances in all datasets are as follows: XLWiC (Figure 5, Figure 6, and Figure 7), MCLWiC (Figures 8, 9, 10, 11, and 12), and AM2iCo (Figures 13, 14, 15, 16, 17, 18, 19, 20, 21, and 22). Similar to § 3, the contextual semantic change-aware axes emerged after the fine-tuning process. Moreover, full results related to the prediction task are as follows: XLWiC (Figure 23), MCLWiC (Figure 24 and Figure 25), AM2iCo (Figure 26, Figure 27, and Figure 28). As shown in § 3, 10% PCA-transformed axes are able to obtain contextual semantic change-aware dimensions.
On the other hand, for the temporal SCD, results for other languages (German, Swedish, and Latin) are shown in Figure 29 and Figure 30. Similar to § 4, temporal semantic change-aware dimensions are observed within 10% PCA-transformed axes. However, there are some difficulties in obtaining these dimensions by PCA-transformed axes with insufficient pretraining data (Swedish) Conneau et al. (2020) or lack of supervision for fine-tuning (Latin) shown in Table 2. In those cases, the use of ICA-transformed axes proved to be effective. More detailed analysis and understanding of those axes for interpretability will be addressed in future work.